Company A is considering a two-stage takeover of Company B. Company B has 2 million shares at current market price of € 20. Company A will offer to buy 51% of Company B shares for € 34 in cash and the remainder with a second bid of 980,000 convertible preferred shares, each of which will be valued at 45% more than B's common stock. A Board of Directors of Company A proposes to make a single offer for all shares of Company B at a price of € 32.50 per share.
Compare the total costs of the two alternatives. Which is preferred in terms of cost minimization?
Answer #1

Alternative 1:

Total Cost = Cash Paid + Value of Preference Shares

= [2000000*0.51*34] + [980000*(20*1.45)]

= 34680000 + 28420000

= 63100000

Alternative 2:

Total Cost = Total Shares*32.5 = 2000000*32.5 = 65000000

Therefore, Total Cost of Alternative 1 i.e. 2 Stage Takeover is LOWER. Hence, It is preferred.

(If this was helpful then please rate positively, Thank You:)